# Rational powers, invariant ideals, and the summation formula

### v2

@@ 2. Rational powers and Rees valuations
The previous lemma yields the following version of the birrational transformation rule for rational powers of ideals (cf. [41, Theorem 9.2.33], [42, Lemme 4.15]).
@@ 3. Rees valuations and rational powers of invariant ideals
In this section we describe the Rees valuations and rational powers of several classes of invariant ideals. This will prove Theorem A. Our approach follows the work on integral closures of these ideals by several authors including [2, 15, 6, 1, 13, 7, 9, 8, 32]. We begin with the following setup.
@@ 3.1. Torus-invariant (monomial) ideals in semigroup rings
Let $\tilde{\mathscr{S}}:=\mathbb{Z}\mathscr{S}\cap\mathbb{R}_{\geqslant 0}\mathscr{S}$ be the normalization of $\mathscr{S}$. By [8, Theorem 4.46] the statement holds for $I\,\mathbb{k}[\tilde{\mathscr{S}}]$. The statement now follows for $I$ by [35, Proposition 1.6.1] since $\mathbb{k}[\mathscr{S}]\subseteq\mathbb{k}[\tilde{\mathscr{S}}]$ is an integral extension [8, Lemma 4.41] and $\Gamma(I\,\mathbb{k}[\tilde{\mathscr{S}}])=\Gamma(I)$. ∎
@@
Clearly $\Gamma(I^{n})=n\Gamma(I)$ for every $n\in{\mathbb{Z}}_{\geqslant 0}$, and so $n\Sigma={\operatorname{conv}}\left(\underline{v}\left(\log(I^{n})\right)\right)$. Also, for every $\mathbf{m}\in{\mathbb{Z}}_{\geqslant 0}^{d}$ the set $\underline{v}^{-1}(\mathbf{m})\cap\mathbf{x}^{\mathscr{S}}$ has at most one element, therefore the assumption in Section missing 3 is trivially satisfied by the non-coordinate supporting hyperplanes of $\Sigma$ The statements then follow from §missing 3.1, Section missing 3, and Section missing 3. ∎
@@
A standard monomial is a product of minors $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{1}\leqslant\dots\leqslant\delta_{p}$. For every $\underline{\delta}$ we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})$ where $\delta_{i}$ is a $s_{i}$-minor and call $\sigma$ the shape of $\underline{\delta}$. Let $\mathcal{B}$ be the set of all standard monomials. By [15, Theorem 2.1] $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For each diagram $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{m}$ we denote by $I_{\sigma}$ the product of determinantal ideals $I_{\sigma}:=I_{s_{1}}\cdots I_{s_{p}}$, and for any subset $\Lambda\subseteq\mathcal{H}_{m}$ we set
@@
For each $t\in[m]$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [15, Theorem 7.1], or [32, Theorem 2.3].
We now show that the ideals $I_{\Lambda}$ have a Rees package. In the following by $I^{(n)}$ we denote the $n$th-symbolic power of the ideal $I$.
@@
By [32, Theorems 2.2 & 2.3] the integral closure $\overline{I_{\Lambda}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{v}(\underline{\delta})\in n\Gamma\}$. By [7, Proposition 7.4.2, Equations (7.2) & (7.3)] the assumption in Section missing 3 is satisfied by the non-coordinate supporting hyperplanes of $\Gamma$; indeed, notice that by [7, Remark 3.2.9] we may assume $\mathbb{k}=\overline{\mathbb{k}}$ and under this assumption Segre products of domains are domains. Thus the statements follow form Section missing 3 and Section missing 3. The last formula follows from [32, Theorems 2.2 & 2.3]. ∎
@@
As defined in [28], the asymptotic resurgence of a homogeneous ideal $I$ in a polynomial ring is $\rho_{a}(I):=\sup\{s/r\mid I^{(st)}\not\subset I^{rt}\text{ for all }t\gg 0\}$. In [23, Theorem 4.10] the authors showed that $\rho_{a}(I)=\max_{v\in{\operatorname{\mathscr{RV}}}(I)}\left\{v(I)/\widehat{v}(I)\right\}$, where $\widehat{v}(I)=\lim_{r\to\infty}v(I^{(r)})/r$. By Theorem 3.2, if $I_{t}$ is the determinantal ideal defined in §missing 3.2, then ${\operatorname{\mathscr{RV}}}(I_{t})=\{\gamma_{1},\ldots,\gamma_{t}\}$. Using [15, Theorem 7.1] and [18, Remark 7.23] one shows that $\rho_{a}(I_{t})=\dfrac{\gamma_{1}(I_{t})}{\displaystyle\widehat{\gamma_{1}}(I_{t})}=\dfrac{t(m-t+1)}{m}$.
For the rest of this subsection assume $\mathbb{k}=\mathbb{C}$ and set $G:=\operatorname{GL}_{m}(\mathbb{C})\times\operatorname{GL}_{n}(\mathbb{C})$. Consider the action of $G$ on $R=\mathbb{C}[X]$ induced by $X_{i,j}\mapsto(AXB^{-1})_{i,j}$ for $(A,B)\in G$. In [15, Theorem 5.1] the authors described certain irreducible representations $L_{\sigma}$ and $\prescript{}{\sigma}{L}$ of $\operatorname{GL}_{m}(\mathbb{C})$ and $\operatorname{GL}_{n}(\mathbb{C})$, respectively, indexed by $\sigma\in\mathcal{H}_{m}$, and they showed that an ideal of $R$ is invariant under the action of $G$ if and only if it is of the form
@@
for some subset $\Lambda\subseteq\mathcal{H}_{m}$. Moreover, in [15, Theorem 8.2] it is shown that $\overline{\mathcal{I}(\Lambda)}=\overline{I_{\Lambda}}$. Thus we obtain the following corollary.
@@
A standard monomial is a product of doset minors $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{1}\leqslant\dots\leqslant\delta_{p}$. For every $\underline{\delta}$ we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})$ where $\delta_{i}$ is a $s_{i}$-minor and call $\sigma$ the shape of $\underline{\delta}$. Let $\mathcal{B}$ be the set of all standard monomials. By [37, §4.2] (see also [2]) $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For each $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{n}$ we denote by $J_{\sigma}$ the product of determinantal ideals $J_{\sigma}:=J_{s_{1}}\cdots J_{s_{p}}$, and for any subset $\Lambda\subseteq\mathcal{H}_{n}$ we set
@@
For each $t\in[n]$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [2, Theorem 5.1], or [32, Theorem 2.7].
@@
By [32, Theorems 2.6 & 2.7] the integral closure $\overline{J_{\Lambda}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{v}(\underline{\delta})\in n\Gamma\}$. By [16, Lemma 5.2] the assumption in Section missing 3 is satisfied by the non-coordinate supporting hyperplanes of $\Gamma$. Thus the statements follow form Section missing 3 and Section missing 3. The last formula follows again from [32, Theorems 2.6 & 2.7]. ∎
@@
for some subset $\Lambda\subseteq\mathcal{H}_{n}$. Moreover, in [2, Theorem 6.1] (see also [32, Theorem 2.11]) it is shown that $\overline{\mathcal{J}(\Lambda)}=\overline{J_{\Lambda}}$. Thus we obtain the following corollary.
@@ 3.4. Sums of products of ideals of Pfaffians and $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals
A standard monomial in this case is a product of Pfaffians $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{1}\leqslant\dots\leqslant\delta_{p}$. For every $\underline{\delta}$ we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})$ where $\delta_{i}$ is a $2s_{i}$-Pfaffian and call $\sigma$ the shape of $\underline{\delta}$. Let $\mathcal{B}$ be the set of all standard monomials. By [37, §4.3] (see also [1]) $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For each $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor n/2\rfloor}$ we denote by $J_{\sigma}$ the product of ideals of Pfaffians $P_{\sigma}:=P_{2s_{1}}\cdots P_{2s_{p}}$, and for any subset $\Lambda\subseteq\mathcal{H}_{\lfloor n/2\rfloor}$ we set
@@
For each $t\in\big{[}\lfloor n/2\rfloor\big{]}$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [1, Theorem 5.1], or [32, Theorem 2.10].
@@
By [32, Theorems 2.9 & 2.10] the integral closure $\overline{P_{\Lambda}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{v}(\underline{\delta})\in n\Gamma\}$. By [16, Lemma 6.2 & proof of Theorem 6.5] and [10, Theorem 1.6] the assumption in Section missing 3 is satisfied by the non-coordinate supporting hyperplanes of $\Gamma$. Thus the statements follow form Section missing 3 and Section missing 3. The last formula follows again from [32, Theorems 2.9 & 2.10]. ∎
@@ 3.5. Products of determinantal ideals of Hankel matrices
For any $t\leqslant\min\{c,n-c+1\}$ let $H_{t}\subset R$ denote the ideal generated by the $t$-minors of $W$. The ideals $H_{t}$ only depends on $t$ and $n$ [13, Corollary 2.2]. Note that the maximum value for $t$ is $\lfloor(n+1)/2\rfloor$. We follow the notation introduced in §missing 3.2 for minors and diagrams.
@@
see [13] for details. The elements in the image $\mathcal{B}$ of $\Phi$ are the standard monomials of $W$. By [13, Theorem 1.2] $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For every standard monomial $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{i}$ a $s_{i}$-maximal-minor and $s_{1}\geqslant\cdots\geqslant s_{p}$, we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor(n+1)/2\rfloor}$. For each $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor(n+1)/2\rfloor}$ we denote by $H_{\sigma}$ the product of determinantal ideals
@@
For each $t\in[n]$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [13, Theorem 3.8].
@@
Adopt §missing 3.5 and the notation above. Let $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor(n+1)/2\rfloor}$. If $\underline{\gamma}:=(\gamma_{1},\ldots,\gamma_{n})$ and $\Gamma:={\operatorname{conv}}\big{(}\underline{\gamma}(\sigma)\big{)}+\mathbb{R}_{\geqslant 0}^{n}$, then $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $H_{\sigma}$. In particular, the Rees valuations of $H_{\sigma}$ are determined by the non-coordinates hyperplanes of $\Gamma$ and ${\operatorname{\mathscr{RV}}}(H_{\sigma})\subseteq\{\gamma_{1},\ldots,\gamma_{s_{1}}\}$. Furthermore, for every $w\in{\mathbb{Q}}_{\geqslant 0}$, $\overline{H_{\sigma}^{w}}$ is spanned as a $\mathbb{k}$-vector space by $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in w\Gamma\}$ and
@@
By [13, Theorem 3.12] the integral closure $\overline{H_{\sigma}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{v}(\underline{\delta})\in n\Gamma\}$. Thus the statements follow form Section missing 3 and Section missing 3. The last formula follows from [13, Theorems 3.8 & 3.12]. ∎

### v3

@@ 2. Rational powers and Rees valuations
The previous lemma yields the following version of the birrational transformation rule for rational powers of ideals (cf. [44, Theorem 9.2.33], [45, Lemme 4.15]).
@@ 3. Rees valuations and rational powers of invariant ideals
In this section we describe the Rees valuations and rational powers of several classes of invariant ideals. This will prove Theorem A. We begin with the following setup.
@@ 3.1. Torus-invariant (monomial) ideals in semigroup rings
Let $\tilde{\mathscr{S}}:=\mathbb{Z}\mathscr{S}\cap\mathbb{R}_{\geqslant 0}\mathscr{S}$ be the normalization of $\mathscr{S}$. By [9, Theorem 4.46] the statement holds for $I\,\mathbb{k}[\tilde{\mathscr{S}}]$. The statement now follows for $I$ by [38, Proposition 1.6.1] since $\mathbb{k}[\mathscr{S}]\subseteq\mathbb{k}[\tilde{\mathscr{S}}]$ is an integral extension [9, Lemma 4.41] and $\Gamma(I\,\mathbb{k}[\tilde{\mathscr{S}}])=\Gamma(I)$. ∎
@@
Clearly $\Gamma(I^{n})=n\Gamma(I)$ for every $n\in{\mathbb{Z}}_{\geqslant 0}$, and so $n\Sigma={\operatorname{conv}}\left(\underline{v}\left(\log(I^{n})\right)\right)$. Also, for every $\mathbf{m}\in{\mathbb{Z}}_{\geqslant 0}^{d}$ the set $\underline{v}^{-1}(\mathbf{m})\cap\mathbf{x}^{\mathscr{S}}$ has at most one element, therefore the assumption in Section missing 3 is trivially satisfied by $\mathbf{x}^{\mathscr{S}}$ and $\underline{v}$. The statements then follow from §missing 3.1, Section missing 3, and Section missing 3. ∎
@@
A standard monomial is a product of minors $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{1}\leqslant\dots\leqslant\delta_{p}$. For every $\underline{\delta}$ we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})\in\mathcal{H}_{m}$ where $\delta_{i}$ is a $s_{i}$-minor and call $\sigma$ the shape of $\underline{\delta}$. Let $\mathcal{B}$ be the set of all standard monomials. By [17, Theorem 2.1] $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For each diagram $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{m}$ we denote by $I_{\sigma}$ the product of determinantal ideals $I_{\sigma}:=I_{s_{1}}\cdots I_{s_{p}}$, and for any subset $\Lambda\subseteq\mathcal{H}_{m}$ we set
@@
For each $t\in[m]$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [17, Theorem 7.1].
We now show that the ideals $I_{\Lambda}$ have a Rees package. In the following, by $I^{(n)}$ we denote the $n$th-symbolic power of the ideal $I$.
@@
By [35, Theorems 2.2 & 2.3] the integral closure $\overline{I_{\Lambda}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in n\Gamma\}$. By [8, Proposition 7.4.2, Equations (7.2) & (7.3)] the assumption in Section missing 3 is satisfied by $\mathcal{B}$ and $\underline{\gamma}$; indeed, notice that by [8, Remark 3.2.9] we may assume $\mathbb{k}=\overline{\mathbb{k}}$, and under this assumption Segre products of domains are domains. Therefore, $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $I_{\Lambda}$ by Section missing 3 and Section missing 3. The last formula follows from [35, Theorems 2.2 & 2.3]. ∎
@@
As defined in [30], the asymptotic resurgence of a homogeneous ideal $I$ in a polynomial ring is $\rho_{a}(I):=\sup\{s/r\mid I^{(st)}\not\subset I^{rt}\text{ for all }t\gg 0\}$. In [26, Theorem 4.10] the authors showed that $\rho_{a}(I)=\max_{v\in{\operatorname{\mathscr{RV}}}(I)}\left\{v(I)/\widehat{v}(I)\right\}$, where $\widehat{v}(I)=\lim_{r\to\infty}v(I^{(r)})/r$. By Theorem 3.2, if $I_{t}$ is the determinantal ideal defined in §missing 3.2, then ${\operatorname{\mathscr{RV}}}(I_{t})=\{\gamma_{1},\ldots,\gamma_{t}\}$. Using [17, Theorem 7.1] and [22, Remark 7.23] one shows that $\rho_{a}(I_{t})=\dfrac{\gamma_{1}(I_{t})}{\displaystyle\widehat{\gamma_{1}}(I_{t})}=\dfrac{t(m-t+1)}{m}$.
For the rest of this subsection assume $\mathbb{k}=\mathbb{C}$ and set $G:=\operatorname{GL}_{m}(\mathbb{C})\times\operatorname{GL}_{n}(\mathbb{C})$. Consider the action of $G$ on $R=\mathbb{C}[X]$ induced by $X_{i,j}\mapsto(AXB^{-1})_{i,j}$ for $(A,B)\in G$. In [17, Theorem 5.1] the authors described certain irreducible representations $L_{\sigma}$ and $\prescript{}{\sigma}{L}$ of $\operatorname{GL}_{m}(\mathbb{C})$ and $\operatorname{GL}_{n}(\mathbb{C})$, respectively, indexed by $\sigma\in\mathcal{H}_{m}$, and they showed that an ideal of $R$ is invariant under the action of $G$ if and only if it is of the form
@@
for some subset $\Lambda\subseteq\mathcal{H}_{m}$. Moreover, in [17, Theorem 8.2] it is shown that $\overline{\mathcal{I}(\Lambda)}=\overline{I_{\Lambda}}$. Thus we obtain the following corollary.
@@
A standard monomial is a product of doset minors $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{1}\leqslant\dots\leqslant\delta_{p}$. For every $\underline{\delta}$ we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})\in\mathcal{H}_{n}$ where $\delta_{i}$ is a $s_{i}$-minor and call $\sigma$ the shape of $\underline{\delta}$. Let $\mathcal{B}$ be the set of all standard monomials. By [40, §4.2] (see also [2]) $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For each $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{n}$ we denote by $J_{\sigma}$ the product of determinantal ideals $J_{\sigma}:=J_{s_{1}}\cdots J_{s_{p}}$, and for any subset $\Lambda\subseteq\mathcal{H}_{n}$ we set
@@
For each $t\in[n]$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [2, Theorem 5.1], or [35, Theorem 2.7].
[…]
Adopt §missing 3.3 and the notation above. Set $\underline{\gamma}:=(\gamma_{1},\ldots,\gamma_{n})$. For every $\mathbf{n}\in{\mathbb{Z}}_{\geqslant 0}^{n}$ let $V_{\mathbf{n}}$ (resp. $V_{>\mathbf{n}}$) be the $\mathbb{k}$-subspace of $R$ spanned by the standard monomials $\underline{\delta}$ such that $\underline{\gamma}(\underline{\delta})=\mathbf{n}$ (resp. $\underline{\gamma}(\underline{\delta})>\mathbf{n}$ componentwise). Then the associated graded ring $\mathcal{G}=\oplus_{\mathbf{n}\in{\mathbb{Z}}_{\geqslant 0}^{n}}V_{\mathbf{n}}/V_{>\mathbf{n}}$ is a domain.
Proof.
Let $X$ be a generic $n\times n$ matrix and $S$ be the subalgebra of $\mathbb{k}[X]$ generated by the products of minors $[1,\ldots,t_{1}|i_{1},\ldots,i_{t_{1}}][1,\ldots,t_{2}|i_{1},\ldots,i_{t_{2}}]$ for some $1\leqslant i_{1},i_{2}\leqslant n$. One proceeds as in [11, proof of Thorem 1.6] via [18, Lemma 5.2 and proof of Lemma 5.3] to show $\mathcal{G}\cong S$. The conclusion follows. ∎
@@
By [35, Theorems 2.6 & 2.7] the integral closure $\overline{J_{\Lambda}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in n\Gamma\}$. By §missing 3.3 the assumption in Section missing 3 is satisfied by $\mathcal{B}$ and $\underline{\gamma}$. Therefore, $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $J_{\Lambda}$ by Section missing 3 and Section missing 3. The last formula follows again from [35, Theorems 2.6 & 2.7]. ∎
@@
for some subset $\Lambda\subseteq\mathcal{H}_{n}$. Moreover, in [2, Theorem 6.1] (see also [35, Theorem 2.11]) it is shown that $\overline{\mathcal{J}(\Lambda)}=\overline{J_{\Lambda}}$. Thus we obtain the following corollary.
@@ 3.4. Sums of products of ideals of Pfaffians and $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals
A standard monomial in this case is a product of Pfaffians $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{1}\leqslant\dots\leqslant\delta_{p}$. For every $\underline{\delta}$ we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor n/2\rfloor}$ where $\delta_{i}$ is a $2s_{i}$-Pfaffian and call $\sigma$ the shape of $\underline{\delta}$. Let $\mathcal{B}$ be the set of all standard monomials. By [40, §4.3] (see also [1]) $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For each $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor n/2\rfloor}$ we denote by $J_{\sigma}$ the product of ideals of Pfaffians $P_{\sigma}:=P_{2s_{1}}\cdots P_{2s_{p}}$, and for any subset $\Lambda\subseteq\mathcal{H}_{\lfloor n/2\rfloor}$ we set
@@
For each $t\in\big{[}\lfloor n/2\rfloor\big{]}$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [1, Theorem 5.1], or [35, Theorem 2.10].
@@
By [35, Theorems 2.9 & 2.10] the integral closure $\overline{P_{\Lambda}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in n\Gamma\}$. By [11, Theorem 1.6 and its proof] (see [18, Lemma 6.2 & proof of Theorem 6.5] to remove the assumption on the characteristic) the assumption in Section missing 3 is satisfied by $\mathcal{B}$ and $\underline{\gamma}$. Therefore, $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $P_{\Lambda}$ by Section missing 3 and Section missing 3. The last formula follows again from [35, Theorems 2.9 & 2.10]. ∎
@@ 3.5. Products of determinantal ideals of Hankel matrices
For any $t\leqslant\min\{c,n-c+1\}$ let $H_{t}\subset R$ denote the ideal generated by the $t$-minors of $W$. The ideals $H_{t}$ only depends on $t$ and $n$ [15, Corollary 2.2]. Note that the maximum value for $t$ is $\lfloor(n+1)/2\rfloor$. We follow the notation introduced in §missing 3.2 for minors and diagrams.
@@
see [15] for details. The elements in the image $\mathcal{B}$ of $\Phi$ are the standard monomials of $W$. By [15, Theorem 1.2] $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For every standard monomial $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{i}$ a $s_{i}$-maximal-minor and $s_{1}\geqslant\cdots\geqslant s_{p}$, we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor(n+1)/2\rfloor}$. For each $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor(n+1)/2\rfloor}$ we denote by $H_{\sigma}$ the product of determinantal ideals
@@
For each $t\in[n]$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [15, Theorem 3.8].
@@
Adopt §missing 3.5 and the notation above. Let $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor(n+1)/2\rfloor}$. If $\underline{\gamma}:=(\gamma_{1},\ldots,\gamma_{n})$ and $\Gamma:=\underline{\gamma}(\sigma)+\mathbb{R}_{\geqslant 0}^{n}$, then $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $H_{\sigma}$. In particular, the Rees valuations of $H_{\sigma}$ are determined by the non-coordinates hyperplanes of $\Gamma$ and ${\operatorname{\mathscr{RV}}}(H_{\sigma})\subseteq\{\gamma_{1},\ldots,\gamma_{s_{1}}\}$. Furthermore, for every $w\in{\mathbb{Q}}_{\geqslant 0}$, $\overline{H_{\sigma}^{w}}$ is spanned as a $\mathbb{k}$-vector space by $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in w\Gamma\}$ and
@@
By [15, Theorem 3.12] the integral closure $\overline{H_{\sigma}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in n\Gamma\}$. Thus the statements follow form Section missing 3 and Section missing 3. The last formula follows from [15, Theorems 3.8 & 3.12]. ∎